OpenStudy (anonymous):

Find the given higher-order derivative. MEDAL WILL BE REWARDED! f''(x)=2-(2/x), f'''(x)

hartnn (hartnn):

just differentiate 2-2/x once

OpenStudy (anonymous):

So wait...let me calculate this...

OpenStudy (anonymous):

Im confused on what to do

hartnn (hartnn):

yeah, \(\dfrac{d}{dx}f''(x)=f'''(x)\)

hartnn (hartnn):

\(\large f'''(x)=\dfrac{d}{dx}(2-\dfrac{2}{x})\)

OpenStudy (anonymous):

so would it be 2^2-x^2?

hartnn (hartnn):

2^2 ?

hartnn (hartnn):

whats d/dx 2 = .. ?

hartnn (hartnn):

derivative of a constant is = ... ?

OpenStudy (anonymous):

it equals -2/dx^3

hartnn (hartnn):

no... derivative of a constant is 0

OpenStudy (anonymous):

So we have 0-2/x?

hartnn (hartnn):

derivative of (1/x) will be ... ?

OpenStudy (anonymous):

-1/x^2

hartnn (hartnn):

yes, so d/dx (-2/x) will be 2/x^2 right ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

thats your f'''(x)

OpenStudy (anonymous):

ohhhhh okay, I see how you go that! Thank you

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

More to come!